An aqueous solution of MgCl2 that is 27.5% (mass%) and has a solution density of 1.25 g/mL. MgCl2 has a molar mass of 95.211 g/mol. What is the molarity of this MgCl2 solution?
Answer #1

Let volume of solution be 1 L

volume , V = 1 L

= 1*10^3 mL

density, d = 1.25 g/mL

use:

mass = density * volume

= 1.25 g/mL *1*10^3 mL

= 1.25*10^3 g

This is mass of solution

mass of MgCl2 = 27.5 % of mass of solution

= 27.5*1250.0/100

= 343.75 g

Molar mass of MgCl2 = 95.211 g/mol

mass(MgCl2)= 343.75 g

use:

number of mol of MgCl2,

n = mass of MgCl2/molar mass of MgCl2

=(3.438*10^2 g)/(95.211 g/mol)

= 3.61 mol

volume , V = 1 L

use:

Molarity,

M = number of mol / volume in L

= 3.61/1

= 3.61 M

Answer: 3.61 M